Preliminary Agenda ================== 1. VISTA general status and planned activities; outcome from the intervention 17-20.06.2010 (TSz); relevant (if any) PPRS summary (VDI) 2. Status of various software: 2.1. SADT 2.2. OT 2.3. IP 3. QC (i.e. persistance), status of the reprocessing (WHu, JLe): in case of telephone connection problems WHu sent the follwoing: I recently distributed the report on the persistence, which collects all important information (I hope). According to Steffens comments, I have translated the power low fit in ADU/sec units. Before, in the report the results were given in ADU/300sec units. I have also corrected the persistence values of the second saturation for the underlying persistence caused by the first saturation which happened one hour earlier. It looks like persistence is not additive. I will discuss the results with PSO, whn I'm back. 3. Validity of twilight flats, triggered by Wolfgang's finding from June 25. When should OBs be re-classified if no proper flatfielding can be performed? (MHi) 4. Progress of the surveys (MRe) 5. Survey progeress, simulations and feedback from the PIs (MRe, VDI) 6. Action Items review (enclosed bellow is the AI section from the minutes of the last IOT) 7. Others 8. Date of the next meeting (end of July, 2010) ************************************************************ ************************************************************ Review of action items from previous IOT meeting (2010-04-23): Progress report on previous action items: * AI 2009-01: Monitor execution times & survey completion status is ---------- being followed up by MAr, MRe. 2009-07-29: there is a first prototype 2009-10-06: no news (MRe) 2009-12-01: no news 2010-03-23: the progress was reported earlier (see above; MRe) let's have another review on the next IOT when we will have more OBs executed (MAr) 2010-04-23: no news X-Original-STATUS: ONGOING [update 2010-05-27] CLOSED, is a permanent issue * AI 2009-11: TSz should find out if the constraints sets could be carried with the OBs, to the OS and the fits header of the VIRCAM files. Otherwise it is not easy to propagate these data by interfaces. 2009-07-29: TSz iterating with R. Schmutzer, but DFI (T. Bierwirth) will also need to be involved. For surveys, given the small range of programs it is expected that there are not too many different OBs, and this might need to be followed in a different/manual way in the beginning - this could imply a major change for the tools that cannot be done in the last moment. 2010-03-23: the keywords have been implemented except for the moon; there is a ticket with details; some testing done by VDI, no errors found; keep ongoing, try to complete during my next shift end of Mar - early Apr. 2010-04-23: VDI reports one formula is missing X-Original-STATUS: ONGOING [update 2010-05-27] Corresponding PPRS is still pending. * AI 2009-14: GHu to check that the seeing distribution function in the scheduler for VIRCAM takes into account the finite (0.51") nominal resolution of the instrument. 2010-03-23: GHu reported that it is not clear if TaToo will be used for scheduling time on VISTA; decided to keep this AI open until it becomes clear how VISTA time will be scheduled. 2010-04-23 no news X-Original-STATUS: ONGOING [update 2010-05-27] Ongoing. Thomas Szeifert states that for the scheduling the seeing distribution is _not_ considered in the requested band. Marina Rejkuba notes that there may be some chance that this is updated by ESO. * AI 2010-01: TSz the calibration database should be updated with new flats and master calibrations and how to do it regularly, as no procedure is currently in place. 2010-01-26: WHu should give Thomas Bierwirth a USB disk with the updated calibrations before Feb7, 2010. Then on try to define a procedure for a regular update of reference frames on Paranal 2010-03-23: done in Feb; the frequency of updates should be three months, the next one will be in May; SMi will talk to DA's to find out how to clean the database, i.e. removing the old calibrations with a script? Once this is done, we close this AI. 2010-04-23: given the intervention, it is not necessary to update the calibDB now with pre-intervention calibrations. X-Original-STATUS: ONGOING [update 2010-05-27] Closed. New DB prepared by WHu (May 21). * AI 2010-02: TSz: how many standard stars fields should be taken? Effort in February to take up to 3 fields. Request to implement a change of jitter pattern by JLe, from 4 to 5 jitter. Try to keep calibrations plan withing 1 hrs total execution time. 2010-01-26: JLe suggests to discuss issues with PIs, and provide feedbacks to TSz, especially in the ZY band calibrations in particular. 2010-03-23: MRe had a discussion with P. Lucas (VVV co-PI) about the frequency of the standards, the feedback is pending, we will keep this open until we hear from him. VDI points that we will overcalibrate at the beginning but at some point we won't need to bother anymore with standards at all because we will have the entire sky calibrated in the NIR (unless one needs better accuracy or in special cases of high extinction regions). A short discussion about feeding this back to the PIs - at the moment it is premature (as well as the information about the expected extra year necessary to complete the surveys) because we are still unsure ourselves. 2010-04-23: no news (MRe not attending) X-Original-STATUS: ONGOING [update 2010-05-27] Closed. Situation cannot be improved with the current set of standards. Corresponding discussion: Marina Rejkuba summarises comments from the user committee. The SMP of VVV asked them to take into account additional time for standards since they do heavily extincted fields. Then ESO decided that users are not charged standard star OBs. The system is calibrating itself, so it should be fine in the end. Thomas Szeifert notes that only the VVV asked for additional standards, doing UKIRT in addition to TSF. They do one UKIRT per night, on one chip only. This needs extrapolation to other chips. He also notes that a 'Landolt-type' project with several FTE would be needed to provide a consistent set of standards with large colour range in the 'non-standard' filters. Mike Irwin notes that it is up to the VVV team to make suggestions on how to optimize their photometric calibration. They can do precise internal, relative, VISTA calibration, but no absolute calibration. * * New Action Items from 2010-03-23: AI 2010-03: completion rate of the surveys to be compared with the new simulation of scheduling which should now include real submitted OBs. This comparison should be used to review the ranking and decide if it is necessary to revise it for the future. MRe to provide completion rate until May 1. GHu+VDI to run a new simulation with realistic OB database. Deadline for this AI is May IOT. 2010-04-23: no news [update 2010-05-27] Ongoing, to be redone including technical time loss. AI 2010-04: TSz to modify the AO priority/validity 2010-04-23: no news (TSz not attending) [update 2010-05-27] Tolerance changed. Closed. AI 2010-05: JLe to incorporate the EXTNAME in the CASU products 2010-04-23: no news (JLe not attending) [update 2010-05-27] Done. closed. AI 2010-06: TSz+VDI to improve the TIO training, to prepare an operational manual 2010-04-23: done X-Original-STATUS: ONGOING [update 2010-05-27] Operations wiki has been updated. More training needed. Ongoing. AI 2010-07: JEm to fix the SADT - 2010-04-23: JEm will fix the most critical items on the bug list, but does not guarantee to fix items that go beyond the original requirements. Some critical items are to be discussed, like the control of parameter values that might be manually edited by the user. The PA bug was identified and will be fixed. The new catalog issue on UCAC-3 isn't solved yet. The tool will be delivered by the end of next week. Ancillary material (documentation) will be distributed. [update 2010-05-27] Ongoing. New corrected version of SADT released, testing by ESO staff until June 8. Summary of corresponding discussion: Jim Emerson notes that tile2 is still not treated properly. It is commonly decided to remove this from the IP for P86. SADT also now reads UCAC3. He states that the remaining issue with proper motion are an issue of athe catalogue, not SADT. He furthermore suggests to spend a night of engineering time to check well what are the correct suitable magnitude ranges for Technical / Guiding chip. Keeping the same allowed magnitude range at a different catalogue (UCAC3) may not necessarily help. Thomas Szeifert adds that there can be very high magnitude uncertainties in the catalogues. For Ultravista the problem is that possible guide stars are generally on the brightness limit. Marina Rejkuba inquires whether the UCAC3 will help Ultravista. Jim Emerson replies that there are less stars in UCAC3 than in GSC. Jim Emerson also asks to send comments/objections to the modifications done to SADT, which are summarised in his recent e-mail. Michael Hilker ask Valentin Ivanov to do SADT tests on the sky during the next week or so. He will try to do this, together with the IP test. Michael Hilker asks for comments regarding SADT until June 8. AI 2010-08: VDI to take care of the new IP by May 2010-04-23: will be done X-Original-STATUS: ONGOING [update 2010-05-27] Ongoing. IP prepared, test is still pending in combination with new SADT. ** New Action Items from meeting 2010-04-23: AI 2010-09: photometric zeropoints as calculated by the pipeline. [update 2010-05-27] Ongoing, see corresponding discussion this meeting AI 2010-10 WHu: It should be investigated, if the quartile of the ellipticity can be used as a threshold for OB rejection. [update 2010-05-27] Ongoing, to be investigated after new de-blending option AI 2010-11 LdB: use a new default value for the de-blending option of the pipeline [update 2010-05-27] Done [note added in proof: done on May 31].